OpenStudy (anonymous):

dy/dx (1/a)arcsec(u/a) u=function a=constant

OpenStudy (anonymous):

\[\Large f(x)=\frac{ 1 }{ a }\sec ^{-1}\frac{ u }{ a }\] Find f'(x)

OpenStudy (anonymous):

u is the function and a is a constant btw

OpenStudy (anonymous):

Yeah, ln(u)u' but I'm trying to work backwards from a formula

OpenStudy (anonymous):

ln(x)

OpenStudy (anonymous):

Hold on

OpenStudy (mathmale):

\[\frac{ d }{ dx }\frac{ 1 }{ a }\sec ^{-1}\frac{ u }{ a }\]

OpenStudy (mathmale):

Hint: 1/a is just a constant coefficient. What could we do with it to simplify the differentiation?

OpenStudy (mathmale):

If that's unclear, start by finding \[\frac{ d }{ dx }\sec ^{-1}x~ for~ practice.\]

OpenStudy (mathmale):

If you're unsure, see the tables in http://www.math.com/tables/derivatives/tableof.htm

OpenStudy (anonymous):

I'm trying to get to this formula by differentiating \[\Large \int\limits\limits_{}^{}\frac{1}{u}\frac{ du }{\sqrt{u^2-a^2} }=\frac{1}{a}arcsec\frac{\left| u \right|}{a}+C\]

OpenStudy (mathmale):

What would happen were you to differentiate the left side of the equation immediately above?

OpenStudy (mathmale):

As for the right side: start out with the simpler \[y=arcsec x\]with the goal of finding the derivative of y with respect to x.

OpenStudy (mathmale):

If y = arcsec x, then x=sec y. Taking the derivative of both sides, \[\frac{ dx }{ dx }=\frac{ d }{ dx }\sec y\]

OpenStudy (mathmale):

Can you finish this differentiation? Hint: use the chain rule last.

OpenStudy (anonymous):

1=y'/abs (y)*(y^2-1)

OpenStudy (anonymous):

1/abs (y)*(y^2-1)=y' And now plug y back in?

OpenStudy (mathmale):

that's right...solve for (dy/dx). I'd bet that your (dy/dx) is similar to the left side of your original equation. Think about what we have done so far. What more remains to do? Note that I used x=cos y as an example; now that you have correctly found (dy/dx), you can apply the same technique to find the derivative of your function with argument (u/a). OK?

OpenStudy (anonymous):

Ok but what I'm stuck with is plugging my argument in for y. I get 1/abs (abs (x)/a)*sqrt(absx)/a) the x ' are supposed to be U ' but yeah I get stuck

OpenStudy (mathmale):

Just a sec, please. I'm going to write this out on paper for myself before trying to explain further what needs to be done.

OpenStudy (anonymous):

I have to go ill come ba c k to this

OpenStudy (mathmale):

Let's see if the following makes sense to you: Given y=arcsec x, x-sec y. Goal: Find (dy/dx). \[\frac{ d }{ dx }[x=\sec y]\rightarrow \frac{ dx }{ dx }=\sec y \tan y \frac{ dy }{dx }\]

OpenStudy (mathmale):

Algebraic manipulation results in \[1=\sec y \tan y \frac{ dy}{ dx}\]

OpenStudy (mathmale):

or\[\frac{ dy }{ dx }=\frac{ 1 }{ \sec y \tan y }\]

OpenStudy (mathmale):

Review this work. What would you do next? We're almost "there."

OpenStudy (anonymous):

Sub the y back in?

OpenStudy (anonymous):

I got it now